OpenStudy (anonymous):

I am forgetting something .0386=x/12, 10=x^-1 or 12=x^2 how do I find x?

OpenStudy (anonymous):

on the first one, .0386 will be multiplied by 12. so that, (.0386)*12=x

OpenStudy (anonymous):

on the second one, 10=1\x x=1\10

OpenStudy (anonymous):

on the third one, find the square rt of both sides.. that makes, sqrt.of 12 = sqrt of x^2 sqrt of 4.3 = x 2sqrtof 3 = x

OpenStudy (anonymous):

remember also that in the third one, you're solving, 12 = x^2, which is a quadratic so there are actually two solutions for this one...

OpenStudy (anonymous):

Thanks, it is always late night math with basics that I doubt my ability to do the basics

OpenStudy (anonymous):

np

OpenStudy (anonymous):

oh yes... so that would be + and - 2sqrt 0f 3... ^_^